OpenStudy (anonymous):

Why is the size of the IP header limited to 60 bytes?

OpenStudy (anonymous):

Internet Header Length (IHL) The second field (4 bits) is the Internet Header Length (IHL), which is the number of 32-bit words in the header. Since an IPv4 header may contain a variable number of options, this field specifies the size of the header (this also coincides with the offset to the data). The minimum value for this field is 5 (RFC 791), which is a length of 5×32 = 160 bits = 20 bytes. Being a 4-bit value, the maximum length is 15 words (15×32 bits) or 480 bits = 60 bytes

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/IPv4

OpenStudy (anonymous):

Thank you for reply. I couldn't understand why is the max # 15. Now i see that it is 4 bit and max # can only be 15 because max 4 bit number is 1111 in Base2= 8 + 4 + 2 +1 In Base10= 15 in Base16. Thank you